OpenStudy (anonymous):

The equation of a line in the xy-plane with slope -1 passing through the point with coordinates (x,y)=(3,0) is

OpenStudy (anonymous):

\(y-y_1=m(x-x_1)\) put \(m=-1, x_1=3, y_1=0\)

OpenStudy (anonymous):

are you still having problems ?

OpenStudy (anonymous):

satellite pretty much just explained how to solve this using the point slope formula. Let me get you started... y - y1 = m(x - x1) using slope(m) -1 and points (3,0) y - 0 = -1(x - 3).....don't forget to distribute the -1 through the parenthesis. If you are still having problems, check this out.... http://www.purplemath.com/modules/strtlneq2.htm

OpenStudy (anonymous):

whats the answer

OpenStudy (anonymous):

are you even gonna try to do this ?

OpenStudy (anonymous):

lets do this together.... y - 0 = -1(x - 3) can you distribute the -1 through the parenthesis ?

OpenStudy (anonymous):

the answer is C usually

OpenStudy (anonymous):

your messing with my brain satellite...lol

OpenStudy (anonymous):

can you distribute the -1 through the parenthesis, @jahvoan ?

OpenStudy (anonymous):

y=-x+3

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

could it be x-y=3

OpenStudy (anonymous):

BINGO......you got it :) y = -x + 3 or in standard form it is x + y = 3

OpenStudy (anonymous):

is x-y=3 wrong?

OpenStudy (anonymous):

yes.....y = -x + 3 if you are going to put this in standard form, and move the x to the other side, then you have to change signs on the x. Let me show you... y = -x + 3 (add x to both sides) y + x = -x + x = 3 (combine like terms) y + x = 3 Do you understand how I did that ?

OpenStudy (anonymous):

as you can see, the x's on the right side cancel each other out.

OpenStudy (anonymous):

your answer is : In slope intercept form y = -x + 3 In standard form y + x = 3 I do not know which form your teacher is asking for,but both of these answers are correct.